OpenStudy (anonymous):

Silly Question (Hard Question) Ok, it's time for Quick Medal Time! Who ever post the correct answer wins (made this one a little harder than −8q + 10 + 1 − 3q) Question: 100% + 3/8 = ?

OpenStudy (anonymous):

you mean 100+3/8 which is = Exact Form: 803/8 decimal form=100.375

OpenStudy (anonymous):

the first one is = −11(q−1)

OpenStudy (anonymous):

@chaser71 Incorrect, answer is 1.375 the "−8q + 10 + 1 − 3q" is not one of the questions in the current Quick Medal Time....

OpenStudy (anonymous):

Since you're the only one who answer the medal goes to you :3

OpenStudy (anonymous):

Its definelty what i put trust me!!!

OpenStudy (anonymous):

decimal form=100.375 thats the decimal form, you never said change the decimal over certain places.